A Bibliography of Publications of Jochen Heinzmann

Jochen Heinzmann Systems Engineering

Research School of Information Sciences and Engineering

RSISE Bld, 115 Australian National University

Canberra, ACT 0200 Australia

Tel: +61 2 6279 8822 FAX: ?n/a?

E-mail: Jochen.Heinzmann@syseng.anu.edu.au (Internet)

05 January 2013 Version 1.02

Abstract

This bibliography records publications of Jochen Heinzmann.

Title word cross-reference

3-D [HZ98a].

= [HMKZ98].

communication [ZH98]. Computer [ZH96b]. Controller [LLH95, LLH96]. Cooperating [LLH95, LLH96].

D [HZ98a]. Distributed [LLH95, LLH96]. Dynamic [LLH95, LLH96].

Estimation [HZ98a, JHZ98].

Face [Hei96, HZ97a]. Facial [HZ98a, ZH96a, ZH96b]. Friendly [HMKZ98].

Gaze [HZ98a]. Gesture [Hei96, HZ97a, ZH96a]. Gestures [ZH96b].

Human [Hei96, HZ97b, HMKZ98, HZ98b, ZH96a, ZH96b, ZH98]. Human-Robot [HZ97b, HZ98b, ZH96a, ZH98].

Interaction [HZ97b, HZ98b, ZH96a, ZH96b]. Interface [HZ97b, HZ98b, ZH98]. Interfaces [HMKZ98].

Mapping [LLH95, LLH96]. Mechanisms [HMKZ98]. Mobile [JHZ98].

novel [ZH98].

Paradigm [HZ98a]. Point [HZ98a]. Pose [HZ98a, JHZ98].

Range [JHZ98]. Real [Hei96, HZ97a, HZ98a, LLH95, LLH96, ZH96b]. Real-Time [Hei96, HZ97a, HZ98a, LLH95, LLH96, ZH96b]. Recognition [Hei96, HZ97a, ZH96a, ZH96b]. Robot [HZ97b, HZ98b, JHZ98, LLH95, LLH96, ZH96a, ZH98]. Robots [HMKZ98]. Robust [HZ97a, HZ98a].

Safe [HMKZ98]. Servoing [JHZ98]. Smart [HMKZ98]. Systems [LLH95, LLH96].

Task [LLH95, LLH96]. Time [Hei96, HZ97a, HZ98a, LLH95, LLH96, ZH96b]. Tracking [Hei96, HZ97a, HZ98a].

using [HZ98a, ZH96a].

Visual [HZ97b, HZ98b, JHZ98, ZH96b, ZH98].

References

Heinzmann:1996:RTH

[Hei96] Jochen Heinzmann. Real-time human face tracking and gesture recognition. Master’s thesis, Universität Karlsruhe, Fakultät für Informatik, 1996. URL

http://wwwsyseng.anu.edu.au/
~jochen/pubs/masterthesis.ps.

Heinzmann:1998:SIS

[HMKZ98] Jochen Heinzmann, Yoshio Matsumoto, Jon Kieffer, and Alexander Zelinsky. Smart interfaces + safe mechanisms = human friendly robots. In IARP International Workshop on Humanoid and Human Friendly Robotics, pages II–3, October 1998. URL

http://wwwsyseng.anu.edu.au/
~jochen/pubs/hfr98.ps.

Heinzmann:1997:RRT

[HZ97a] Jochen Heinzmann and Alexander Zelinsky. Robust real-time face tracking and gesture recognition. In Proceedings of the International Joint Conference on Artificial Intelligence, IJCAI’97, volume 2, pages 1525–1530, August 1997.

URL http://wwwsyseng.anu.
edu.au/~jochen/pubs/ijcai97.
ps.

Heinzmann:1997:VIH

[HZ97b] Jochen Heinzmann and Alexander Zelinsky. A visual interface for human-robot interaction. In Proceedings of the International Conference on Field and Service Robotics, FSR’97, pages 572–579, December 1997. URL

http://wwwsyseng.anu.edu.au/
~jochen/pubs/fsr97.ps.

Heinzmann:1998:DFP

[HZ98a] Jochen Heinzmann and Alexander Zelinsky. 3-D facial pose and gaze point estimation using a robust real-time tracking paradigm. In Proceedings of the IEEE International Conference on Automatic Face and Gesture Recognition, pages 142–147, April 1998. URL

http://wwwsyseng.anu.edu.au/
~jochen/pubs/fg98.ps.

Heinzmann:1998:VIH

[HZ98b] Jochen Heinzmann and Alexander Zelinsky. A visual interface for human-robot interaction. In Zelinsky A., editor, Field and Service Robotics, pages 540–547. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1998. URL

http://wwwsyseng.anu.edu.au/
~jochen/pubs/fsr97.ps.

Jung:1998:RPE

[JHZ98] David Jung, Jochen Heinzmann, and Alexander Zelinsky. Range and pose estimation for visual servoing of a mobile robot. In Proceedings of the IEEE International Conference on Robotics and Automation, pages 1226–1231, May 1998. URL

http://wwwsyseng.anu.edu.au/
~jochen/pubs/icra98.ps.

Lueth:1995:DTM

[LLH95] Tim Lüth, Thomas Längle, and Jochen Heinzmann. Dynamic task mapping in a real-time controller for distributed cooperating robot systems. In 13th IFAC Workshop on Distributed Computer Control Systems, DCCS’95, 1995.

Lueth:1996:DTM

[LLH96] Tim Lüth, Thomas Längle, and Jochen Heinzmann. Dynamic task mapping in a real-time controller for distributed cooperating robot systems. Control Engineering Practice, 4(6):831–838, 1996.

Zelinsky:1996:HRI

[ZH96a] Alexander Zelinsky and Jochen Heinzmann. Human-robot interaction using facial gesture recognition. In Proceedings of the International Workshop on Robot and Human Communication, pages 256– 261. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, November 1996.

Zelinsky:1996:RTV

[ZH96b] Alexander Zelinsky and Jochen Heinzmann. Real-time visual recognition of facial gestures for human computer interaction. In Proceedings of the International Conference on Automatic Face and Gesture Recognition, pages 351–356. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1996.

                                        URL http://wwwsyseng.anu.
                                        edu.au/~jochen/pubs/fg96.ps.

Zelinsky:1998:NVI

[ZH98] Alexander Zelinsky and Jochen Heinzmann. A novel visual interface for human-robot communication. Advanced Robotics, 11(8): 827–852, 1998.